Jane took three hours to drive to her friend's home on the weekend. Her average speed was 85km/hr for the first two hours. What was her average speed for the last hour of the trip if her average speed was 75km/hr for the whole trip?
Answer by ptaylor(2048)

Distance(d) equals Rate(r) times Time(t) or d=rt; r=d/t and t=d/r
also
Average Speed =(Total Distance)/(Total Time)
Total Distance in this problem is distance traveled for first 2 hours plus distance traveled in last hour
Total Time =3 hours
For the first two hours, distance traveled was 85*2=170 mi
Let d=distance traveled in the last hour
Soooo, our equation to solve is:
75=(170+d)/3
225=170+d
d=55 miles----distance traveled in last hour
Average Speed for last hour was 55/1=55 mph
CK
75=225/3
75=75
Hope this helps---ptaylor